OpenStudy (anonymous):

*** WILL MEDAL *** An electrician earns $110 after his first hour of working for a client. His total pay based on the number of hours worked can be represented using the sequence shown. 110, 130, 150, 170, ... Which recursive formula can be used to determine the total amount of money earned for each successive hour worked based on the amount of money currently earned? f(n + 1) = f(n) + 20 f(n + 1) = f(n) + 110 f(n + 1) = f(n + 1) + 20 f(n + 1) = f(n + 1) + 110

OpenStudy (cmtboy2):

f(n + 1) = f(n) + 110 i think

OpenStudy (mathstudent55):

The first two numbers in the sequence are 110 and 130. Let's say that f(n), for some value of n, is 110. The next n, n + 1, has f(n+1) = 130. What is the difference between f(n+1) and f(n)? What is the difference between 130 and 110?